OpenStudy (linn99123):

A cone has a base radius of 7 centimeters and a height of 21 centimeters. What is the volume of the cone? A. 28π cm3 B. 49π cm3 C. 147π cm3 D. 343π cm3 @undeadknight26

OpenStudy (linn99123):

i think its A

undeadknight26 (undeadknight26):

i got 2 formulas...would u know which one was correct? i got: 1/3 × pi × r^2 × h and 1/3 × b × h

undeadknight26 (undeadknight26):

either formula i use i get a weird answer...with the first formula(1/3 × pi × r^2 × h)i got 609 and with the second formula i got 48.5...

OpenStudy (linn99123):

it would be B then

OpenStudy (accessdenied):

Generally figures which come to a single point (the tip of a cone, or a pyramid) have 1/3 the area of their prism counterpart. Both formulas are correctly written as: 1/3 x Area of base x height But a cone has a circular base, so volume = 1/3 x (pi r^2) x height

OpenStudy (accessdenied):

radius is 7, and height was 21: Plug it all in: 1/3 * pi (7)^2 * 21 And leave pi as it is. :)

OpenStudy (linn99123):

how do i do the ^ part :o

OpenStudy (accessdenied):

7^2 is an exponent exponent just means repeated multiplying. So 7^2 means multiply 7 two times, or 7 times 7.

OpenStudy (accessdenied):

Similarly 3^3 would mean multiply 3 three times, or 3 * 3 * 3. Does that make sense?

OpenStudy (linn99123):

yes

OpenStudy (accessdenied):

So: 1/3 * pi 7^2 * 21 1/3 * 7 * 7 * 21 * pi <-- that pi is just going to come along for the ride now. We could use a calculator, or notice that 21 = 7 * 3. The 3 and 1/3 multiply together to make 1, so we have: 1/3 * 7 * 7 * (3 * 7) * pi (1/3 * 3) * 7 * 7 * 7 * pi 7 * 7 * 7 * pi

OpenStudy (accessdenied):

Its not common to have the cubed numbers memorized, so at this point we can use a calculator again: 7 * 7 * 7 = 343

OpenStudy (linn99123):

tyhat makes alot of sense :o